# Wireless Communication for Low-Altitude Economy with UAV Swarm Enabled Two-Level Movable Antenna System

### v1

@@ III-B1 Optimization of ${\bf V}$ With Given ${\bf Q}$
$\mathop{\max}\limits_{\_}{\bf{V}}$ $\ \mathop{\min}\limits_{\_}{k\in{\cal K}}\ {R_{\_}k}$ (23)
${\rm{s.t.}}$ $\ \left\|{{{\bf{v}}_{\_}k}}\right\|=1,\ \forall k.$
A closer look at (8) shows that the receive beamforming ${{\bf{v}}_{\_}k}$ only impacts the SINR for UE $k$. Thus, problem (23) can be decomposed as $K$ parallel sub-problems. Specifically, the sub-problem for UE $k$ is written as
$\mathop{\max}\limits_{\_}{{\bf{v}}_{\_}k}$ $\ \ {\gamma_{\_}k}$ (24)
${\rm{s.t.}}$ $\ \left\|{{{\bf{v}}_{\_}k}}\right\|=1.$
Since the resulting SINR in (8) is a generalized Rayleigh quotient w.r.t. ${\bf{v}}_{\_}k$, the optimal ${\bf{v}}_{\_}k$ to (24) is given by the classical MMSE beamforming [46]:
${\bf{v}}_{\_}k^{\star}=\frac{{{\bf{C}}_{\_}k^{-1}{{\bf{h}}_{\_}k}}}{{\left\|{{\bf{C}}_{\_}k^{-1}{{\bf{h}}_{\_}k}}\right\|}},\ \forall n,k.$ (25)
III-B2 Optimization of ${{\bf{q}}_{\_}l}$ With Given ${\bf V}$ and $\left\{{{\bf{q}}_{\_}{l^{\prime}},\forall l^{\prime}\neq l}\right\}$
For given ${\bf V}$ and $\left\{{{\bf{q}}_{\_}{l^{\prime}},\forall l^{\prime}\neq l}\right\}$, the sub-problem of (11) for optimizing the placement position of UAV $l$ can be expressed as
$\mathop{\max}\limits_{\_}{{\bf{q}}_{\_}l}$ $\ \mathop{\min}\limits_{\_}{k\in{\cal K}}{R_{\_}k}$ (26)
${\rm{s.t.}}$ $\ {{\bf{q}}_{\_}l}\in{\cal C},$
${\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|^{2}}\geq d_{\_}{\min}^{2},\ \forall l^{\prime}\neq l.$
By introducing the slack variable $\rho$, problem (26) is equivalently written as
$\mathop{\max}\limits_{\_}{{\bf{q}}_{\_}l,\rho}$ $\ \rho$ (27)
${\rm{s.t.}}$ $\ {R_{\_}k}\geq\rho,\ \forall k\in{\cal K},$ (27a)
$\ {{\bf{q}}_{\_}l}\in{\cal C},$ (27b)
$\ {\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|^{2}}\geq d_{\_}{\min}^{2},\ \forall l^{\prime}\neq l.$ (27c)
Moreover, to tackle the non-convexity of the constraint (27a), the slack variables $\left\{{{\eta_{\_}k},{\mu_{\_}k},\forall k}\right\}$ are introduced such that
${e^{{\eta_{\_}k}}}=1+\sum\limits_{\_}{i=1}^{K}{{{\bar{P}}_{\_}i}{{\left|{{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}i}}\right|}^{2}}},$ (28)
${e^{{\mu_{\_}k}}}=1+\sum\limits_{\_}{i=1,i\neq k}^{K}{{{\bar{P}}_{\_}i}{{\left|{{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}i}}\right|}^{2}}}.$ (29)
@@
$\mathop{\max}\limits_{\_}{{{\bf q}_{\_}l},\rho,\left\{{{\eta_{\_}k},{\mu_{\_}k}}\right\}_{\_}{k=1}^{K}}\ \ \rho$ (30)
${\rm{s.t.}}$ $\ {\frac{{\ln\left({{e^{{\eta_{\_}k}-{\mu_{\_}k}}}}\right)}}{{\ln 2}}}\geq\rho,\ \forall k\in{\cal K},$ (30a)
${\sigma^{2}}+\sum\limits_{\_}{i=1}^{K}{{P_{\_}i}{{\left|{{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}i}}\right|}^{2}}}\geq{\sigma^{2}}{e^{{\eta_{\_}k}}},\ \forall k\in{\cal K},$ (30b)
${\sigma^{2}}+\sum\limits_{\_}{i=1,i\neq k}^{K}{{P_{\_}i}{{\left|{{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}i}}\right|}^{2}}}\leq{\sigma^{2}}{e^{{\mu_{\_}k}}},\ \forall k\in{\cal K},$ (30c)
${{\bf{q}}_{\_}l}\in{\cal C},$ (30d)
${\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|^{2}}\geq d_{\_}{\min}^{2},\ l^{\prime}\neq l.$ (30e)
@@
Note that problem (30) is still challenging to be solved since the constraints (30b), (30c) and (30e) are non-convex. To tackle this issue, the SCA technique [3] is applied in the following.
@@
Let ${{\bf{\bar{h}}}_{\_}{i,l}}\in{{\mathbb{C}}^{\left({L-1}\right)\times 1}}$ denote the resulting channel after removing ${h_{\_}{i,l}}$ from ${{\bf{h}}_{\_}i}$, and ${{\bf{\bar{v}}}_{\_}{k,l}}\in{{\mathbb{C}}^{\left({L-1}\right)\times 1}}$ denote the resulting beamforming vector after removing ${v_{\_}{k,l}}$ from ${{\bf{v}}_{\_}k}$, where ${h_{\_}{i,l}}$ and ${v_{\_}{k,l}}$ denote the $l$-th element of ${{\bf{h}}_{\_}i}$ and ${{\bf{v}}_{\_}k}$, respectively. Then, ${f_{\_}{k,i}}\triangleq{\left|{{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}i}}\right|^{2}}$ can be expressed as
${f_{\_}{k,i}}$ $={\bf{h}}_{\_}i^{H}{{\bf{v}}_{\_}k}{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}i}=\underbrace{\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{\mathop{\rm Re}\nolimits}\left\{{h_{\_}{i,l}^{*}{V_{\_}{k,l,l^{\prime}}}{h_{\_}{i,l^{\prime}}}}\right\}}}_{\_}{{g_{\_}{k,i,l}}}$ (31)
$+\underbrace{{{\left|{{\alpha_{\_}i}}\right|}^{2}}{V_{\_}{k,l,l}}+{\bf{\bar{h}}}_{\_}{i,l}^{H}{{\bf{\bar{V}}}_{\_}{k,l,l}}{{\bf{\bar{h}}}_{\_}{i,l}}}_{\_}{{{\bar{g}}_{\_}{k,i,l}}},$
@@
where ${V_{\_}{k,l,l^{\prime}}}\triangleq{v_{\_}{k,l}}v_{\_}{k,l^{\prime}}^{*}$ and ${{\bf{\bar{V}}}_{\_}{k,l,l}}\in{{\mathbb{C}}^{\left({L-1}\right)\times\left({L-1}\right)}}\triangleq{{\bf{\bar{v}}}_{\_}{k,l}}{\bf{\bar{v}}}_{\_}{k,l}^{H}$, and ${\bar{g}}_{\_}{k,i,l}$ is independent of ${{\bf{q}}_{\_}l}$. By substituting ${h_{\_}{i,l}}={\alpha_{\_}i}{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{\_}i}\left({{{\bf{q}}_{\_}l}}\right)}}$ into ${g_{\_}{k,i,l}}$, we have
${g_{\_}{k,i,l}}=\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}{\rm{Re}}\left\{{{V_{\_}{k,l,l^{\prime}}}{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}\left({{\varphi_{\_}i}\left({{{\bf{q}}_{\_}l}}\right)-{\varphi_{\_}i}\left({{{\bf{q}}_{\_}{l^{\prime}}}}\right)}\right)}}}\right\}}$ (32)
@@
=_$=\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|}\cos\left({\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)-\angle{V_{\_}{k,l,l^{\prime}}}}\right).$
@@
It is observed that (32) is neither convex nor concave with respect to ${{\bf q}_{\_}l}$. To this end, two quadratic surrogate functions are respectively constructed to serve as the lower and upper bounds for ${g_{\_}{k,i,l}}$, shown in the following lemma.
@@
The lower and upper bounds of ${g_{\_}{k,i,l}}$ are respectively given by
${g_{\_}{k,i,l}}$ $\geq{\left[{{g_{\_}{k,i,l}}}\right]_{\_}{\rm lb}}={g_{\_}{k,i,l}}\left({{\bf{q}}_{\_}l^{\left(j\right)}}\right)+\nabla{g_{\_}{k,i,l}}{\left({{\bf{q}}_{\_}l^{\left(j\right)}}\right)^{T}}\left({{{\bf{q}}_{\_}l}-{\bf{q}}_{\_}l^{\left(j\right)}}\right)$ (33)
$\ \ \ \ \ \ \ \ -\frac{{{\delta_{\_}{k,i,l}}}}{2}{\left({{{\bf{q}}_{\_}l}-{\bf{q}}_{\_}l^{\left(j\right)}}\right)^{T}}\left({{{\bf{q}}_{\_}l}-{\bf{q}}_{\_}l^{\left(j\right)}}\right),$
${g_{\_}{k,i,l}}$ $\leq{\left[{{g_{\_}{k,i,l}}}\right]_{\_}{\rm ub}}={g_{\_}{k,i,l}}\left({{\bf{q}}_{\_}l^{\left(j\right)}}\right)+\nabla{g_{\_}{k,i,l}}{\left({{\bf{q}}_{\_}l^{\left(j\right)}}\right)^{T}}\left({{{\bf{q}}_{\_}l}-{\bf{q}}_{\_}l^{\left(j\right)}}\right)$ (34)
$\ \ \ \ \ \ \ \ +\frac{{{\delta_{\_}{k,i,l}}}}{2}{\left({{{\bf{q}}_{\_}l}-{\bf{q}}_{\_}l^{\left(j\right)}}\right)^{T}}\left({{{\bf{q}}_{\_}l}-{\bf{q}}_{\_}l^{\left(j\right)}}\right),$
where ${\delta_{\_}{k,i,l}}=\frac{{8{\pi^{2}}}}{{{\lambda^{2}}}}{\left|{{\alpha_{\_}i}}\right|^{2}}\sum\nolimits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{\left|{V_{\_}{k,l,l^{\prime}}}\right|}$, ${{\bf{q}}_{\_}l^{\left(j\right)}}$ denotes the resulting placement position of UAV $l$ in the $j$-th iteration, and $\nabla{g_{\_}{k,i,l}}({{\bf{q}}_{\_}l^{\left(j\right)}})$ denotes the gradient of ${{g_{\_}{k,i,l}}}$ over ${\bf{q}}_{\_}l^{\left(j\right)}$.
@@
With Lemma 1, the lower and upper bounds of the function ${f_{\_}{k,i}}$ are
${f_{\_}{k,i}}\geq{\left[{{f_{\_}{k,i}}}\right]_{\_}{\rm lb}}\triangleq{\left[{{g_{\_}{k,i,l}}}\right]_{\_}{\rm lb}}+{{\bar{g}}_{\_}{k,i,l}},$ (35)
${f_{\_}{k,i}}\leq{\left[{{f_{\_}{k,i}}}\right]_{\_}{\rm ub}}\triangleq{\left[{{g_{\_}{k,i,l}}}\right]_{\_}{\rm ub}}+{{\bar{g}}_{\_}{k,i,l}},$ (36)
which are concave and convex with respect to ${{\bf{q}}_{\_}l}$, respectively.
Moreover, to tackle the non-convexity of ${e^{{\mu_{\_}k}}}$ on the right-hand-side (RHS) of (30c), let $\mu_{\_}k^{\left(j\right)}$ denote the resulting variable in the $j$-th iteration; then the convex function ${e^{{\mu_{\_}k}}}$ is lower-bounded by
${e^{{\mu_{\_}k}}}\geq{\left[{{e^{{\mu_{\_}k}}}}\right]_{\_}{{\rm{lb}}}}\triangleq{\mu_{\_}k}{e^{\mu_{\_}k^{\left(j\right)}}}+\left({1-\mu_{\_}k^{\left(j\right)}}\right){e^{\mu_{\_}k^{\left(j\right)}}}.$ (37)
Similarly, for given ${\bf{q}}_{\_}l^{\left(j\right)}$, the convex function ${\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|^{2}}$ on the left-hand-side (LHS) of (30e) is lower-bounded by [3]
${\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|^{2}}\geq{\left[{{{\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|}^{2}}}\right]_{\_}{{\rm{lb}}}}$ (38)
$\triangleq 2{\left({{\bf{q}}_{\_}l^{\left(j\right)}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)^{T}}\left({{{\bf{q}}_{\_}l}-{\bf{q}}_{\_}l^{\left(j\right)}}\right)+{\left\|{{\bf{q}}_{\_}l^{\left(j\right)}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|^{2}}.$
As a result, problem (30) is lower-bounded by the following problem for given $\{{{\bf{q}}_{\_}l^{\left(j\right)},\mu_{\_}k^{\left(j\right)}}\}$.
$\mathop{\max}\limits_{\_}{{{\bf{q}}_{\_}l},\rho,\left\{{{\eta_{\_}k},{\mu_{\_}k}}\right\}_{\_}{k=1}^{K}}\ \ \rho$ (39)
${\rm{s.t.}}$ $\ {{\eta_{\_}k}-{\mu_{\_}k}}\geq\rho\ln 2,\ \forall k\in{\cal K},$
${\sigma^{2}}+\sum\limits_{\_}{i=1}^{K}{{P_{\_}i}{{\left[{{f_{\_}{ki}}}\right]}_{\_}{{\rm{lb}}}}}\geq{\sigma^{2}}{e^{{\eta_{\_}k}}},\ \forall k\in{\cal K},$
${\sigma^{2}}+\sum\limits_{\_}{i=1,i\neq k}^{K}{{P_{\_}i}{{\left[{{f_{\_}{ki}}}\right]}_{\_}{{\rm{ub}}}}}\leq{\sigma^{2}}{\left[{{e^{{\mu_{\_}k}}}}\right]_{\_}{{\rm{lb}}}},\ \forall k\in{\cal K},$
${{\bf{q}}_{\_}l}\in{\cal C},$
${\left[{{{\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|}^{2}}}\right]_{\_}{{\rm{lb}}}}\geq d_{\_}{\min}^{2},\ \forall l^{\prime}\neq l,$
@@
The overall algorithm is summarized in Algorithm 1. It is worth mentioning that the objective value at each iteration is non-decreasing, thus guaranteeing the convergence of Algorithm 1. Moreover, the computational complexity is analyzed as follows. In step 3, the complexity for obtaining the receive beamforming is ${\cal O}({K{L^{3}}})$. The complexity from step 4 to step 6 is approximately ${\cal O}({I_{\_}1}L{{({2K})}^{3}})$, where $I_{\_}1$ denotes the maximum number of iterations required by SCA for convergence. Thus, the total computational complexity of Algorithm 1 is ${\cal O}({{I_{\_}2}K{L^{3}}+{I_{\_}1}{I_{\_}2}L{({2K})}^{3}})$, where $I_{\_}2$ denotes the number of iterations required by the alternating optimization for convergence.
@@
1:  Initialize ${{\bf{Q}}^{\left(j\right)}}$ and $\{{{\bf{v}}_{\_}k^{\left(j\right)}}\}$ randomly, and let $j=0$.
@@
3:     For given ${{\bf{Q}}^{\left(j\right)}}$, obtain the optimal solution to (24) based on (25), denoted as $\{{{\bf{v}}_{\_}k^{\left({j+1}\right)}}\}$.
@@
5:      Obtain $\{{\bf{q}}_{\_}l^{\left({j+1}\right)}\}$ by solving problem (39), given  $\{{\bf{q}}_{\_}1^{\left({j+1}\right)},\cdots,{\bf{q}}_{\_}{l-1}^{\left({j+1}\right)},{\bf{q}}_{\_}{l+1}^{\left(j\right)},\cdots,{\bf{q}}_{\_}L^{\left(j\right)}\}$ and $\{{{\bf{v}}_{\_}k^{\left({j+1}\right)}}\}$.

### v2

@@ III-B1 Optimization of $\bf V$} With Given $\bf Q$}
$\mathop{\max}\limits_{\bf{V},\rho}$ $\ \rho$ (24)
${\rm{s.t.}}$ $\ {R_{k}}\geq\rho,\ \forall k,$
$\ \left\|{{{\bf{v}}_{k}}}\right\|=1,\ \forall k.$
A closer look at (9) shows that the receive beamforming ${\bf{v}}_{k}$} only impacts the SINR for UE $k$, and (9) is a generalized Rayleigh quotient w.r.t. ${\bf{v}}_{k}$, the optimal ${\bf{v}}_{k}$ to (24) is given by the classical MMSE beamforming [47]:
${\bf{v}}_{k}^{\star}=\frac{{{\bf{C}}_{k}^{-1}{{\bf{h}}_{k}}}}{{\left\|{{\bf{C}}_{k}^{-1}{{\bf{h}}_{k}}}\right\|}},\ \forall k.$ (25)
III-B2 Optimization of ${\bf{q}}_{l}$} With Given $\bf V$} and $\left\{{{\bf{q}}_{l^{\prime}},\forall l^{\prime}\neq l}\right\}$
For given $\bf V$} and $\left\{{{\bf{q}}_{l^{\prime}},\forall l^{\prime}\neq l}\right\}$, the sub-problem of (12) for optimizing the placement position of UAV $l$ can be expressed as
$\mathop{\max}\limits_{{\bf{q}}_{l},\rho}$ $\ \rho$ (26)
${\rm{s.t.}}$ $\ {R_{k}}\geq\rho,\ \forall k\in{\cal K},$ (26a)
$\ {{\bf{q}}_{r}}+{{\bf{q}}_{l}}\in{\cal C},$ (26b)
$\ {\left\|{{{\bf{q}}_{l}}-{{\bf{q}}_{l^{\prime}}}}\right\|^{2}}\geq d_{\min}^{2},\ \forall l^{\prime}\neq l.$ (26c)
Moreover, to tackle the non-convexity of the constraint (26a), the slack variables $\left\{{{\eta_{k}},{\mu_{k}},\forall k}\right\}$ are introduced such that
${e^{{\eta_{k}}}}=1+\sum\limits_{i=1}^{K}{{{\bar{P}}_{i}}{{\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|}^{2}}},$ (27)
${e^{{\mu_{k}}}}=1+\sum\limits_{i=1,i\neq k}^{K}{{{\bar{P}}_{i}}{{\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|}^{2}}}.$ (28)
@@
$\mathop{\max}\limits_{{{\bf q}_{l}},\rho,\left\{{{\eta_{k}},{\mu_{k}}}\right\}_{k=1}^{K}}\ \ \rho$ (29)
${\rm{s.t.}}$ $\ {\frac{{\ln\left({{e^{{\eta_{k}}-{\mu_{k}}}}}\right)}}{{\ln 2}}}\geq\rho,\ \forall k\in{\cal K},$ (29a)
${\sigma^{2}}+\sum\limits_{i=1}^{K}{{P_{i}}{{\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|}^{2}}}\geq{\sigma^{2}}{e^{{\eta_{k}}}},\ \forall k\in{\cal K},$ (29b)
${\sigma^{2}}+\sum\limits_{i=1,i\neq k}^{K}{{P_{i}}{{\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|}^{2}}}\leq{\sigma^{2}}{e^{{\mu_{k}}}},\ \forall k\in{\cal K},$ (29c)
${{\bf{q}}_{r}}+{{\bf{q}}_{l}}\in{\cal C},$ (29d)
${\left\|{{{\bf{q}}_{l}}-{{\bf{q}}_{l^{\prime}}}}\right\|^{2}}\geq d_{\min}^{2},\ l^{\prime}\neq l.$ (29e)
@@
Note that problem (29) is still challenging to be solved since the constraints (29b), (29c) and (29e) are non-convex. To tackle this issue, the SCA technique [3] is applied in the following, which is an efficient iterative optimization technique that successively updates the optimization variables by solving the approximated convex problem at each iteration.
@@
Let ${{\bf{\bar{h}}}_{i,l}}\in{{\mathbb{C}}^{\left({L-1}\right)\times 1}}$ denote the resulting channel after removing $h_{i,l}$} from ${\bf{h}}_{i}$}, and ${{\bf{\bar{v}}}_{k,l}}\in{{\mathbb{C}}^{\left({L-1}\right)\times 1}}$ denote the resulting beamforming vector after removing $v_{k,l}$} from ${\bf{v}}_{k}$}, where $h_{i,l}$} and $v_{k,l}$} denote the $l$-th element of ${\bf{h}}_{i}$} and ${\bf{v}}_{k}$}, respectively. Then, ${f_{k,i}}\triangleq{\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|^{2}}$ can be expressed as
${f_{k,i}}$ $={\bf{h}}_{i}^{H}{{\bf{v}}_{k}}{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}=\underbrace{\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{\mathop{\rm Re}\nolimits}\left\{{h_{i,l}^{*}{V_{k,l,l^{\prime}}}{h_{i,l^{\prime}}}}\right\}}}_{{g_{k,i,l}}}$ (30)
$+\underbrace{{{\left|{{\alpha_{i}}}\right|}^{2}}{V_{k,l,l}}+{\bf{\bar{h}}}_{i,l}^{H}{{\bf{\bar{V}}}_{k,l,l}}{{\bf{\bar{h}}}_{i,l}}}_{{{\bar{g}}_{k,i,l}}},$
@@
where ${V_{k,l,l^{\prime}}}\triangleq{v_{k,l}}v_{k,l^{\prime}}^{*}$, ${{\bf{\bar{V}}}_{k,l,l}}\in{{\mathbb{C}}^{\left({L-1}\right)\times\left({L-1}\right)}}\triangleq{{\bf{\bar{v}}}_{k,l}}{\bf{\bar{v}}}_{k,l}^{H}$, and ${\bar{g}}_{k,i,l}$ is independent of ${\bf{q}}_{l}$}. By substituting ${h_{i,l}}={\alpha_{i}}{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{r_{i,l}}}}$ into $g_{k,i,l}$}, we have
${g_{k,i,l}}=\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}{\rm{Re}}\left\{{{V_{k,l,l^{\prime}}}{e^{{\rm{j}}\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)}}}\right\}}$ (31)
@@
$=\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}\cos\left({\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)+\angle{V_{k,l,l^{\prime}}}}\right).$
@@
It is observed that (31) is neither convex nor concave w.r.t. ${\bf q}_{l}$}. To this end, two quadratic surrogate functions are respectively constructed to serve as the lower and upper bounds for $g_{k,i,l}$}, shown in the following lemma.
@@
The lower and upper bounds of $g_{k,i,l}$} are respectively given by
${g_{k,i,l}}$ $\geq{\left[{{g_{k,i,l}}}\right]_{\rm lb}}={g_{k,i,l}}\left({{\bf{q}}_{l}^{\left(j\right)}}\right)+\nabla{g_{k,i,l}}{\left({{\bf{q}}_{l}^{\left(j\right)}}\right)^{T}}\left({{{\bf{q}}_{l}}-{\bf{q}}_{l}^{\left(j\right)}}\right)$ (32)
$\ \ \ \ \ \ \ \ -\frac{{{\delta_{k,i,l}}}}{2}{\left({{{\bf{q}}_{l}}-{\bf{q}}_{l}^{\left(j\right)}}\right)^{T}}\left({{{\bf{q}}_{l}}-{\bf{q}}_{l}^{\left(j\right)}}\right),$
${g_{k,i,l}}$ $\leq{\left[{{g_{k,i,l}}}\right]_{\rm ub}}={g_{k,i,l}}\left({{\bf{q}}_{l}^{\left(j\right)}}\right)+\nabla{g_{k,i,l}}{\left({{\bf{q}}_{l}^{\left(j\right)}}\right)^{T}}\left({{{\bf{q}}_{l}}-{\bf{q}}_{l}^{\left(j\right)}}\right)$ (33)
$\ \ \ \ \ \ \ \ +\frac{{{\delta_{k,i,l}}}}{2}{\left({{{\bf{q}}_{l}}-{\bf{q}}_{l}^{\left(j\right)}}\right)^{T}}\left({{{\bf{q}}_{l}}-{\bf{q}}_{l}^{\left(j\right)}}\right),$
where ${\delta_{k,i,l}}=\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}\sqrt{2+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{{C_{\max}}}}{{r_{i}^{2}}}}\right)}^{2}}}$, we have ${\nabla^{2}}{g_{k,i,l}}\preceq{\delta_{k,i,l}}{\bf{I}}$, ${\bf{q}}_{l}^{\left(j\right)}$} denotes the resulting placement position of UAV $l$ in the $j$-th iteration, and $\nabla{g_{k,i,l}}({{\bf{q}}_{l}^{\left(j\right)}})$ denotes the gradient of $g_{k,i,l}$}} over ${\bf{q}}_{l}^{\left(j\right)}$.
@@
With Lemma 1, the lower and upper bounds of the function $f_{k,i}$} are
${f_{k,i}}\geq{\left[{{f_{k,i}}}\right]_{\rm lb}}\triangleq{\left[{{g_{k,i,l}}}\right]_{\rm lb}}+{{\bar{g}}_{k,i,l}},$ (34)
${f_{k,i}}\leq{\left[{{f_{k,i}}}\right]_{\rm ub}}\triangleq{\left[{{g_{k,i,l}}}\right]_{\rm ub}}+{{\bar{g}}_{k,i,l}},$ (35)
which are concave and convex w.r.t. ${\bf{q}}_{l}$}, respectively.
Moreover, to tackle the non-convexity of $e^{{\mu_{k}}}$} on the right-hand-side (RHS) of (29c), let $\mu_{k}^{\left(j\right)}$ denote the resulting variable in the $j$-th iteration; then the convex function $e^{{\mu_{k}}}$} is lower-bounded by
${e^{{\mu_{k}}}}\geq{\left[{{e^{{\mu_{k}}}}}\right]_{{\rm{lb}}}}\triangleq{\mu_{k}}{e^{\mu_{k}^{\left(j\right)}}}+\left({1-\mu_{k}^{\left(j\right)}}\right){e^{\mu_{k}^{\left(j\right)}}}.$ (36)
Similarly, for given ${\bf{q}}_{l}^{\left(j\right)}$, the convex function $\left\|{{{\bf{q}}_{l}}-{{\bf{q}}_{l^{\prime}}}}\right\|^{2}$} on the left-hand-side (LHS) of (29e) is lower-bounded by [3]
${\left\|{{{\bf{q}}_{l}}-{{\bf{q}}_{l^{\prime}}}}\right\|^{2}}\geq{\left[{{{\left\|{{{\bf{q}}_{l}}-{{\bf{q}}_{l^{\prime}}}}\right\|}^{2}}}\right]_{{\rm{lb}}}}$ (37)
$\triangleq 2{\left({{\bf{q}}_{l}^{\left(j\right)}-{{\bf{q}}_{l^{\prime}}}}\right)^{T}}\left({{{\bf{q}}_{l}}-{\bf{q}}_{l}^{\left(j\right)}}\right)+{\left\|{{\bf{q}}_{l}^{\left(j\right)}-{{\bf{q}}_{l^{\prime}}}}\right\|^{2}}.$
As a result, problem (29) is lower-bounded by the following problem for given $\{{{\bf{q}}_{l}^{\left(j\right)},\mu_{k}^{\left(j\right)}}\}$.
$\mathop{\max}\limits_{{{\bf{q}}_{l}},\rho,\left\{{{\eta_{k}},{\mu_{k}}}\right\}_{k=1}^{K}}\ \ \rho$ (38)
${\rm{s.t.}}$ $\ {{\eta_{k}}-{\mu_{k}}}\geq\rho\ln 2,\ \forall k\in{\cal K},$
${\sigma^{2}}+\sum\limits_{i=1}^{K}{{P_{i}}{{\left[{{f_{ki}}}\right]}_{{\rm{lb}}}}}\geq{\sigma^{2}}{e^{{\eta_{k}}}},\ \forall k\in{\cal K},$
${\sigma^{2}}+\sum\limits_{i=1,i\neq k}^{K}{{P_{i}}{{\left[{{f_{ki}}}\right]}_{{\rm{ub}}}}}\leq{\sigma^{2}}{\left[{{e^{{\mu_{k}}}}}\right]_{{\rm{lb}}}},\ \forall k\in{\cal K},$
${{\bf{q}}_{r}}+{{\bf{q}}_{l}}\in{\cal C},$
${\left[{{{\left\|{{{\bf{q}}_{l}}-{{\bf{q}}_{l^{\prime}}}}\right\|}^{2}}}\right]_{{\rm{lb}}}}\geq d_{\min}^{2},\ \forall l^{\prime}\neq l,$
@@
The overall algorithm is summarized in Algorithm 1, and its convergence proof is given as follows.
Proposition 1
The objective value of problem (12) obtained by Algorithm 1 is guaranteed to converge.
Proof:
Please refer to Appendix C. ∎
It is worth mentioning that the objective value at each iteration is non-decreasing, thus guaranteeing the convergence of Algorithm 1. Moreover, the computational complexity is analyzed as follows. In step 3, the complexity for obtaining the receive beamforming is ${\cal O}({K{L^{3}}})$. The complexity from step 4 to step 6 is approximately ${\cal O}({I_{1}}L{{({2K})}^{3}})$, where $I_{1}$ denotes the maximum number of iterations required by SCA for convergence. Thus, the total computational complexity of Algorithm 1 is ${\cal O}({{I_{2}}K{L^{3}}+{I_{1}}{I_{2}}L{({2K})}^{3}})$, where $I_{2}$ denotes the number of iterations required by the alternating optimization for convergence.
@@
1: Initialize ${\bf{Q}}^{\left(0\right)}$} and $\{{{\bf{v}}_{k}^{\left(0\right)}}\}$ randomly, and let $j=0$.
@@
3:  For given ${\bf{Q}}^{\left(j\right)}$}, obtain the optimal solution to (24) based on (25), denoted as $\{{{\bf{v}}_{k}^{\left({j+1}\right)}}\}$.
@@
5:   Obtain $\{{\bf{q}}_{l}^{\left({j+1}\right)}\}$ by solving problem (38), given  $\{{\bf{q}}_{1}^{\left({j+1}\right)},\cdots,{\bf{q}}_{l-1}^{\left({j+1}\right)},{\bf{q}}_{l}^{\left(j\right)},\cdots,{\bf{q}}_{L}^{\left(j\right)}\}$ and $\{{{\bf{v}}_{k}^{\left({j+1}\right)}}\}$.